# SEREP: Semantic Facial Expression Representation for Robust In-the-Wild Capture and Retargeting

### v2

@@ Abstract
Monocular facial performance capture in-the-wild is challenging due to varied capture conditions, face shapes, and expressions. Most current methods rely on linear 3D Morphable Models, which represent facial expressions independently of identity at the vertex displacement level. We propose SEREP (Semantic Expression Representation), a model that disentangles expression from identity at the semantic level. It first learns an expression representation from unpaired 3D facial expressions using a cycle consistency loss. Then we train a model to predict expression from monocular images using a novel semi-supervised scheme that relies on domain adaptation. In addition, we introduce MultiREX, a benchmark addressing the lack of evaluation resources for the expression capture task. Our experiments show that SEREP outperforms state-of-the-art methods, capturing challenging expressions and transferring them to novel identities. Project page is available on the following link.
@@
Figure 1: SEREP captures facial expressions in challenging in-the-wild conditions, including extreme expressions and side views. These are effectively retargeted to a new head geometry preserving the target identity.
@@ 1 Introduction
3D facial performance capture is essential for applications ranging from virtual communication to video games, where users interact with avatars that can be rendered from different viewpoints. While established technologies such as light stages [12] and head-mounted cameras allow for high-fidelity facial capture, they are costly and impractical for general end-users. This motivates research on capturing 3D facial expressions in natural, unconstrained environments using everyday devices like webcams and smartphones.
Capturing 3D facial performance can be viewed as two subtasks: (i) face geometry reconstruction to uncover identity (i.e., face at a neutral expression) and (ii) time-dependent geometry deformation capture to describe facial expressions. Representing these two types of information in disentangled spaces allows for 3D facial reenactment, often called animation retargeting, by applying the acting performance to a character with a different morphology. This implies that facial expressions are represented at a semantic level instead of direct vertex displacements which are coupled with the geometry of a given identity. In addition to allowing for reenactment, disentangled representations can improve performance on downstream tasks [6, 35, 36] while reducing sample complexity [2, 42]. In this paper, we focus on learning a meaningful facial expression representation, assuming that we already have access to a neutral face geometry.
Most models for in-the-wild face expression capture [13, 17, 10, 20, 11, 45, 39] rely on 3D Morphable Models (3DMM) [33, 28] to represent a facial mesh through independent identity and expression bases. In these models, expressions are represented as additive vertex displacements to a neutral mesh. In reality, a semantic facial expression (e.g., wide smile) cannot be simply represented by vertex displacements that can readily be applied to any identity. Mesh deformations inherent to an expression directly depend on the underlying face morphology (i.e., identity). For instance, a smiling expression results in different displacements on two persons, depending on their facial traits, such as the shape of the mouth or the fat tissue structure. This indicates that disentanglement of expression and identity must happen at the semantic level instead of the vertex level as formulated in 3DMMs.
In addition to not truly capturing the semantics of expression, 3DMM-based methods struggle to capture accurate 3D face meshes. While they show promising results when viewed from the original camera angle, issues appear in novel views. This is particularly noticeable in images where the subject does not directly face the camera. In these cases, viewing the 3D reconstruction from other angles reveals artifacts such as asymmetric mouth shapes and distorted expressions. These issues are explained by the self-supervised training strategy which optimizes models using differentiable rendering [29] without explicit 3D supervision. While this approach allows for training on large-scale in-the-wild data, it primarily enforces consistency in the original camera space but produces geometric inconsistencies that are only revealed when changing viewpoints. Some methods [46, 47] circumvent these issues by training exclusively on realistic synthetic data, however modeling and rendering realistic variations in skin, eyes, hair, clothing, and accessories requires significant effort and expertise.
Another key challenge in advancing in-the-wild face capture is the lack of standardized evaluation protocols to measure geometric accuracy. Current evaluation protocols rely on qualitative visual comparisons and on projecting geometry into camera view to use image-space metrics like similarity measures [27], emotion recognition [39, 10] or alignment of face segmentation masks [1, 40]. This reprojection step introduces inaccuracies in measurement and does not allow for evaluating the whole face geometry due to limited points of view. While benchmarks like NOW [41] and REALY [7] exist for evaluating the 3D geometry of neutral expressions, there is no public benchmark for assessing the geometric accuracy of captured facial animations.
In this paper, we propose a method for monocular 3D face capture named SEREP, that addresses the fundamental challenges described above. We introduce a semantic face expression basis, learned from unpaired 3D poses, without supervision in the form of expression, FACS, or emotion labels. We validate the expressivity of our semantic expression representation by measuring the reconstruction error from ground truth expressive 3D meshes and measure how identity is preserved when transferring expression between identities. Our results indicate better identity preservation compared to traditional 3DMM methods.
To address the lack of standardized evaluation protocols measuring geometric accuracy in 3D space, we introduce MultiREX, a new benchmark based on the Multiface dataset [51]. MultiREX contains challenging expressions, different viewpoints of the same subject, and high-quality ground truth geometry, allowing for meaningful reconstruction accuracy evaluation. We show that SEREP is more accurate for 3D performance capture than competing methods and more robust to viewpoints than recent state-of-the-art methods.
Finally, we present visual results for in-the-wild capture and retargeting, which demonstrate improved results on challenging expressions. Overall, our method improves the state-of-the-art in terms of identity preservation after expression transfer, and the geometric reconstruction of expressions from monocular images, especially on non-frontal views. Our contributions are summed up as follows:
(a) Expression capture
(b) Retargeting
Figure 2: Overview of the expression capture and retargeting processes. Given an input image, we extract a semantic expression code $\textbf{z}_{\text{exp}}$ using the encoder $\mathbf{E}_{\text{img}}$. Left: The decoder $\mathbf{D}_{\text{mesh}}$ applies this code to the neutral mesh to obtain the expressive 3D mesh. Right: In a retargeting scenario, the same semantic expression code is used with a target neutral mesh of any other subject.
  • We introduce an identity-aware expression representation that encodes the semantics of facial expression.
  • We propose a semi-supervised training strategy that combines synthetic and in-the-wild data through domain adaptation.
@@
  • We create and publicly release MultiREX, the first video-based geometric facial expression reconstruction benchmark. We see this contribution as an important step towards the advancement of research on face geometry reconstruction from monocular images and videos.
@@ 2 Related work
Facial performance capture from monocular video is an ill-posed problem, as the task is to recover 3D geometry from a 2D projection of the face in the image. To make this problem tractable, most recent methods [17, 58, 10, 20, 11, 45, 31, 8], rely on statistical morphable models (3DMM) [16], such as [33, 28, 32, 8] to constrain the problem. These morphable models use a large collection of scans and define: (i) a template mesh $\overline{\boldsymbol{T}}$ (average face shape), (ii) an identity (shape) basis $\boldsymbol{S}$, and (iii) an expression basis $\boldsymbol{E}$. An arbitrary mesh $M$ is then defined by identity ($\boldsymbol{\beta}$) and expression coefficients ($\boldsymbol{\psi}$) of these linear bases: $M=\overline{\boldsymbol{T}}+\boldsymbol{\beta}\boldsymbol{S}+\boldsymbol{\psi}\boldsymbol{E}$. We note that a given expression vector $\boldsymbol{\psi}$ results in the same vertex displacement regardless of the identity and, therefore, cannot capture the semantic meaning of an expression. In other words, the same semantic expression of two subjects would have different displacements and, therefore, different expression codes.
@@
Some recent work proposed approaches for semantic expression disentanglement. Facescape [53] considers a corpus of multiple identities performing the same set of expressions, and represents identity and expression with a bilinear model for identity and expression [44]. Neural Parametric Head Models [23] also uses a corpus of multiple identities performing the same expressions. This model operates in an implicit representation, where expression deformations are represented with an MLP that takes both expression and identity coefficients as input. We note that both methods require paired data for training (multiple subjects performing the same expression). This has two drawbacks: (i) it is costly to acquire, (ii) for a given expression label, they assume that the subjects are performing exactly the same facial movements. In reality, subjects are not always capable of following the exact instructions for these poses. Other works [54, 9] require facial animations to be decomposed in FACS-based rigs, a task that normally requires a significant amount of manual work. In contrast, we learn a semantic expression space with unpaired data, without any labels or rig decomposition, in a more similar vein to the work done in face reenactment [15, 52].
@@
Most recent methods for in-the-wild face capture train models in a large collection of in-the-wild images, in a self-supervised manner [17, 58, 10, 20, 11, 14, 45]. These methods estimate the identity and expression parameters of a 3DMM, as well as camera, light, and reflectance information of a scene, then use a differentiable renderer [29, 43] to render a synthetic image of the parametrized scene. The model is trained to minimize the difference between rendered images and ground truth, together with auxiliary losses. Since these losses are only based on camera-space information, we observe that these methods can sometimes produce implausible expressions when viewed from other angles, notably when the processed subject is not facing the camera.
@@
Most existing work on monocular face reconstruction provides either (i) qualitative results, (ii) quantitative results on image-centric metrics, or (iii) geometric error in reconstructing neutral meshes. Image-centric metrics include emotion classification [10, 39], and the Intersection-Over-Union of segmented face regions in the predicted mesh, compared to a pseudo-ground truth [1, 45]. Geometric reconstruction error is commonly computed using either the NOW [41] and REALY [7] benchmarks, which only provide data in a neutral pose, being unsuitable for evaluating expression recovery. FaceWarehouse [5] allows for assessing expression capture in static 3D meshes, but provides low-quality scans from RGBD data and does not include multiview capture.
In this work, we propose a new public benchmark for geometric expression capture based on the Multiface [51] dataset. This dataset was collected from a light stage with multiple cameras and provides higher-quality reconstructions. Our proposed benchmark allows for geometric evaluation of expression capture from monocular videos under controlled camera angles.
@@ 3 Method
We propose a two-stage training to capture facial expressions from in-the-wild images and videos. In the first stage, we train a semantic face expression model (Sec. 3.1) that disentangles expression from identity in 3D facial meshes. In the second stage, we train a facial performance capture model (Sec. 3.3) to predict expression codes from monocular images. The final 3D mesh is produced by combining the expression code with a neutral mesh, as illustrated in Fig. 2. The target neutral mesh can be estimated using methods such as [58, 14], and is outside of the scope of this work.
Figure 3: Expression retargeting model. The model learns to reconstruct a source expression $E_{s}$ given the source neutral $N_{s}$ or to retarget it to a target neutral $N_{t}$. The cycle loss aligns the codes $\textbf{z}_{\text{exp}}$ for the same semantic expressions of different subjects.
3.1 Semantic face expression model
We train the semantic face expression model on a retargeting task as illustrated in Fig. 3. Training samples consist of a neutral mesh $N_{s}$, an expressive mesh $E_{s}$ from the same source subject, and a neutral mesh $N_{t}$ from a target subject. The task of the model is to (i) reconstruct the ground truth expression $E_{s}$, and (ii) retarget the expression to the target subject.
This model is composed of an encoder $\mathbf{E}_{\text{id}}$ that encodes identity, an encoder $\mathbf{E}_{\text{exp}}$ that encodes expression, and a single decoder $\mathbf{D}_{\text{mesh}}$ that decodes the final mesh. All models operate directly on mesh vertices, using spiral convolutions [4]. Both encoders consist of 5 blocks containing a spiral convolution followed by a graph pooling operation [37]. The last layer is a fully-connected layer, that produces the identity and expression codes. The decoder consists of a fully-connected layer followed by 5 blocks of graph upsampling and spiral convolutions. We use skip connections between intermediate features of the identity encoder and the mesh decoder. We describe the loss terms used in training below.
@@ 3.1 Semantic face expression model
$\mathcal{L}_{\text{ret}}=L_{\text{rec}}+L_{\text{cycle}}+L_{\text{delta}}+L_{\text{edge}}+L_{\text{eyes}}$ (1)
@@
Figure 4: Face expression capture. Given a synthetic or real input image, the model estimates the expression code $\textbf{z}_{\text{exp}}$, landmarks $\mathbf{l}$, and the domain (real or synthetic). Gradient reversal (GR) and the classifier $\mathbf{C}_{d}$ are used to bring real and synthetic images close in feature space.
@@ 3.2 Synthetic data generation
Once the semantic face expression model is trained, we exploit it to generate a synthetic dataset that will be used to train the face expression capture model. First, we sample a target identity $N_{t}$ and expression code $\textbf{z}_{\text{exp}}$ from the dataset and produce an expressive mesh $\hat{E}_{t}$. We procedurally place a set of teeth and eyes, and a random face albedo texture, environment map, and head pose. Finally, we render a synthetic image using [21]. Unlike [46], we do not add accessories, hair, or clothes to the scene, avoiding any further 3D modeling. Examples of the generated images can be visualized in the supplementary material. We also extract 2D facial landmarks, denoted $\mathbf{l}$, by projecting the corresponding mesh vertices to camera coordinates.
3.3 Face expression capture model
Our face expression capture model estimates a semantic expression code given an image, as illustrated in Fig. 4. The model is jointly trained on three tasks enabling joint training on synthetic and real data. The encoder $\mathbf{E}_{\text{img}}$ extracts latent features $\phi$ from the image. From these shared features, the head $\mathbf{H}_{\text{exp}}$ estimates the expression code $\hat{\textbf{z}}_{\text{exp}}$, and the head $\mathbf{H}_{\text{lmks}}$ estimates the 2D landmarks $\hat{\mathbf{l}}$. A classifier $\mathbf{C}_{d}$ predicts whether the sample is a real or synthetic image for domain adaptation.
@@ 3.3 Face expression capture model
Domain loss. To minimize the domain gap between synthetic and real data, we adopt the domain adaptation approach from [22]. The classifier $\mathbf{C}_{d}$ is trained to distinguish between real and synthetic images. During gradient descent optimization, we reverse the sign of the gradient after $\mathbf{C}_{d}$ and scale its magnitude by a factor $\lambda$. As a result, $\mathbf{C}_{d}$ is optimized to discriminate between the two domains, while the features $\phi$ are encouraged to discard domain-specific information, thus aligning the distributions of both domains more closely.
@@
$\mathcal{L}_{\text{cap}}=L_{\text{code}}+L_{\text{lmks}}+L_{\text{domain}}$ (2)
In summary, the model learns to encode images into a semantic expression code $\textbf{z}_{\text{exp}}$, using the ground truth codes from the synthetic data. The landmark and domain loss terms align the features from real and synthetic images allowing the model to generalize to in-the-wild images. Importance weights in Eq. 2 are omitted for clarity and detailed in the supplementary material.

### v3

@@ Abstract
Monocular facial performance capture in-the-wild is challenging due to varied capture conditions, face shapes, and expressions. Most current methods rely on linear 3D Morphable Models, which represent facial expressions independently of identity at the vertex displacement level. We propose SEREP (Semantic Expression Representation), a model that disentangles expression from identity at the semantic level. We start by learning an expression representation from high-quality 3D data of unpaired facial expressions. Then, we train a model to predict expression from monocular images relying on a novel semi-supervised scheme using low quality synthetic data. In addition, we introduce MultiREX, a benchmark addressing the lack of evaluation resources for the expression capture task. Our experiments show that SEREP outperforms state-of-the-art methods, capturing challenging expressions and transferring them to new identities. Results in the project page.
@@
Figure 1: SEREP extracts facial expressions from in-the-wild monocular images and applies them to a given neutral mesh, enabling accurate expression reconstruction and identity-preserving retargeting.
@@ 1 Introduction
Facial performance capture is essential for applications where users interact with 3D avatars, such as multimedia and video games. For widespread adoption, capturing facial expressions from everyday devices like webcams and smartphones is crucial. Furthermore, for creative applications with fictional characters, the ability to retarget captured expressions to different facial morphologies is equally important, allowing the preservation of performance expressivity while adapting to unique character appearances. Capturing a facial performance that can be applied to different characters necessitates the disentanglement of identity from expression information, which is our motivation for this work.
Current methods for in-the-wild facial expression capture [11, 15, 9, 18, 10, 41, 36] predominantly rely on 3D Morphable Models (3DMM) [31, 26], which represent facial expressions as additive vertex displacements to a neutral mesh. However, this representation has a fundamental limitation: expression-driven deformations inherently depend on identity-specific facial features, such as mouth shape or fat distribution. In 3DMMs, when the same expression parameters are applied to different identities, they produce identical vertex displacements regardless of the target’s facial structure, leading to identity leakage and unnatural results. This suggests that expressions should be represented at the semantic level rather than through direct vertex displacements.
In this paper, we present SEREP (Semantic Expression Representation), a method for monocular 3D face capture that disentangles identity and expression at the semantic level. By semantic level, we mean representing expressions in a feature space that can be transformed into identity-specific vertex displacements, allowing the same underlying expression to manifest differently across various facial geometries. SEREP learns a semantic face expression basis from 3D expressive faces, without requiring paired expression data, FACS, or emotion labels. This expression basis allows to generate diverse synthetic faces with minimal manual effort, which we use to train our face expression capture model through a novel semi-supervised learning scheme. By learning from both low-quality synthetic and in-the-wild data, SEREP performs accurate monocular facial expression capture, that is robust in unconstrained scenarios.
Beyond limitations in expression representation, existing methods face significant training challenges. Some require complex self-supervised approaches with differentiable rendering for image-space reprojection [11, 15, 9, 18, 10, 41, 36], while others require substantial manual effort to create synthetic training data that is as realistic as possible [42, 43]. SEREP is trained directly in 3D space, thus circumvent the challenges associated with rendering model outputs, and generalizes to in-the-wild data, without relying on expensive and complex realistic synthetic data.
As a final contribution, we propose MultiREX, a new benchmark for facial expression capture, addressing the current lack of benchmarks for geometric evaluation of expression capture. Current evaluation protocols typically rely on image-space metrics [11, 15, 1, 37] which can introduce inaccuracies and fail to comprehensively evaluate 3D facial geometry. MultiREX, based on the Multiface dataset [47], features challenging expressions with multiple viewpoints, and allows measuring geometric accuracy in 3D space.
We validate our approach through extensive experiments that show both the expressivity of our semantic expression basis and its ability to preserve identity during retargeting. SEREP outperforms state-of-the-art methods in terms of capture accuracy on MultiREX and in-the-wild data, exhibits greater robustness to viewpoint changes and better preserves identity during expression transfer.
Figure 2: We first train a semantic expression model to disentangle expression from identity on 3D meshes. Next, we generate a synthetic dataset of 2D renders. Finally, we train an expression capture model on synthetic and real images, that estimates an expression code from a monocular image. This code, along with a neutral mesh, is used by $\mathbf{D}_{\text{mesh}}$ to obtain an expressive 3D mesh.
The contributions of our paper are as follows:
  • We introduce an identity-aware decoder that represents expressions at the semantic level, reducing identity leakage during expression transfer.
  • We propose a semi-supervised training scheme using synthetic data that can be generated without substantial manual effort, while still generalizing to in-the-wild conditions.
@@
  • We publicly release MultiREX, the first video-based geometric 3D facial expression reconstruction benchmark.
@@ 2 Related work
Facial performance capture from monocular video is an ill-posed problem, as the task is to recover 3D geometry from a 2D projection of the face in the image. To make this problem tractable, most recent methods [15, 54, 9, 18, 10, 41, 29, 6], rely on statistical morphable models (3DMM) [14], such as [31, 26, 30, 6] to constrain the problem. These morphable models use a large collection of scans and define: (i) a template mesh $\overline{\boldsymbol{T}}$ (average face shape), (ii) an identity (shape) basis $\boldsymbol{S}$, and (iii) an expression basis $\boldsymbol{E}$. An arbitrary mesh $M$ is then defined by identity ($\boldsymbol{\beta}$) and expression ($\boldsymbol{\psi}$) coefficients of these linear bases: $M=\overline{\boldsymbol{T}}+\boldsymbol{\beta}\boldsymbol{S}+\boldsymbol{\psi}\boldsymbol{E}$. We note that a given expression vector $\boldsymbol{\psi}$ results in the same vertex displacement regardless of the identity and, therefore, cannot capture the semantic meaning of an expression. In other words, the same semantic expression of two subjects would have different displacements and, therefore, different expression codes.
@@
Some recent work proposed approaches for semantic expression disentanglement. Facescape [49] considers a corpus of multiple identities performing the same set of expressions, and represents identity and expression with a bilinear model for identity and expression [40]. Neural Parametric Head Models (NPHM) [21] (and per extension MonoNPHM [22]) uses a corpus of multiple identities performing the same expressions. This model operates in an implicit representation, where expression deformations are represented with an MLP that takes both expression and identity coefficients as input. We note that both methods require paired data for training (multiple subjects performing the same expression), as is commonly done in re-targeting papers [7]. This has two drawbacks: (i) it is costly to acquire, (ii) for a given expression label, they assume that the subjects are performing exactly the same facial movements. In reality, subjects are not always capable of following the exact instructions for these poses. Other works [50, 8] require facial animations to be decomposed in FACS-based rigs, a task that normally requires a significant amount of manual work. In contrast, we learn a semantic expression space with unpaired data, without any labels or rig decomposition, in a more similar vein to the work done in face reenactment [13, 48].
@@
Most recent methods for in-the-wild face capture train models in a large collection of in-the-wild images, in a self-supervised manner [15, 54, 9, 18, 10, 12, 41, 32]. These methods estimate the identity and expression parameters of a 3DMM, as well as camera, light, and reflectance information of a scene, then use a differentiable renderer [27, 39] to render a synthetic image of the parametrized scene. The model is trained to minimize the difference between rendered images and ground truth, together with auxiliary losses. This camera-space optimization can sometimes result in implausible expressions when viewed from other angles, notably when the processed subject is not facing the camera.
@@
Most existing work on monocular face reconstruction provides qualitative results along with either (i) geometric error in reconstructing neutral meshes or (ii) quantitative results using image-centric metrics. NOW [38] and REALY [5] benchmarks are commonly used to evaluate neutral mesh geometric reconstruction error, but do not account for expression recovery. Image-centric metrics include emotion classification [9, 36], and the Intersection-Over-Union of segmented face regions in the predicted mesh, compared to a pseudo-ground truth [1, 41]. FaceWarehouse [4] enables 3D expression evaluation instead, but only offers low-quality RGBD scans without multiview capture. In this work, we propose a new public benchmark for geometric expression capture based on the Multiface [47] dataset. This dataset was collected from a light stage with multiple cameras, and provides high-quality reconstructions with corresponding monocular videos under controlled camera angles.
@@ 3 Method
We propose a three-stage approach to capture facial expressions from in-the-wild images (Fig. 2). First, we train a semantic expression model (Sec. 3.1) that disentangles expression from identity in 3D facial meshes. This model’s retargeting ability is leveraged in the second stage (Sec. 3.2) to generate 2D synthetic data to which we know the corresponding expression codes. In the third stage (Sec. 3.3), we train a facial performance capture model to predict expression codes from monocular images. In the absence of target expression codes for in-the-wild data (e.g., no corresponding 3D ground-truth), the model learns to regress expression codes for synthetic data and relies on a domain adaptation strategy with real data to bridge the domain gap.
Figure 3: Semantic expression model training. The model learns to reconstruct a source expression $E_{s}$ given the source neutral $N_{s}$, and to retarget it to a target neutral $N_{t}$. The cycle loss aligns the codes $\textbf{z}_{\text{exp}}$ for the same semantic expressions of different subjects.
3.1 Semantic expression model
The task of the semantic expression model is to represent the expression from an expressive mesh $E_{s}$ and apply it to a neutral mesh of the same source subject $N_{s}$ (i.e., reconstruction) or a different target subject $N_{t}$ (i.e., retargeting). For training, the model learns those two objectives together, as illustrated in Fig. 3.
This model is composed of an encoder $\mathbf{E}_{\text{id}}$ that encodes identity, an encoder $\mathbf{E}_{\text{exp}}$ that encodes expression, and a single decoder $\mathbf{D}_{\text{mesh}}$ that decodes the final mesh. All models operate directly on mesh vertices, using spiral convolutions [3, 34]. We describe the loss terms used in training below, where the first two are designed for learning a semantic expression space, while the remaining serve as regularization losses.
@@ 3.1 Semantic expression model
$\mathcal{L}_{\text{ret}}=L_{\text{rec}}+L_{\text{cycle}}+L_{\text{edge}}+L_{\text{eyes}}+L_{\text{delta}}$ (1)
@@
Figure 4: Expression capture model training. Given a synthetic or real input image, the model encodes features $\phi$ and estimates the expression code $\textbf{z}_{\text{exp}}$, landmarks $\mathbf{l}$, and the domain (real or synthetic). Gradient reversal and the classifier $\mathbf{C}_{d}$ are used to bring real and synthetic images close in feature space.
@@ 3.2 Synthetic data generation
Once the semantic expression model is trained, we exploit it to generate a synthetic dataset that will be used to train the face expression capture model. First, we sample a target identity $N_{t}$ and expression code $\textbf{z}_{\text{exp}}$ from the dataset and produce an expressive mesh $\hat{E}_{t}$. We procedurally place a set of teeth and eyes, and a random face albedo texture, environment map, and head pose. Finally, we render a synthetic image using [19]. Unlike [42], we do not add accessories, hair, or clothes to the scene, avoiding any further 3D modeling. Examples of the generated images can be visualized in the supplementary material. We also extract 2D facial landmarks, denoted $\mathbf{l}$, by projecting the corresponding mesh vertices to camera coordinates.
3.3 Expression capture model
Our expression capture model estimates semantic expression codes from images, using a joint training strategy with synthetic and real data (Fig. 4). The encoder $\mathbf{E}_{\text{img}}$ extracts latent features $\phi$ which feed three task-specific components: the $\mathbf{H}_{\text{exp}}$ head estimates the semantic expression code $\hat{\textbf{z}}_{\text{exp}}$ using the ground truth codes from the synthetic data, $\mathbf{H}_{\text{lmks}}$ estimates 2D landmarks $\hat{\mathbf{l}}$, and $\mathbf{C}_{d}$ classifies images as real or sythetic. The last two are used to align features across domains to improve generalization to in-the-wild images.
@@ 3.3 Expression capture model
Domain loss. To minimize the domain gap between synthetic and real data, we adopt the domain adaptation approach from [20], and define the corresponding domain loss $L_{d}$. The classifier $\mathbf{C}_{d}$ is trained to distinguish between real and synthetic images. During gradient descent optimization, we reverse the sign of the gradient after $\mathbf{C}_{d}$ and scale its magnitude by a factor $\lambda$. As a result, $\mathbf{C}_{d}$ is optimized to discriminate between the two domains, while the features $\phi$ are encouraged to discard domain-specific information, thus aligning the distributions of both domains more closely.
@@
$\mathcal{L}_{\text{cap}}=L_{\text{code}}+L_{\text{lmks}}+L_{\text{d}}$ (2)
Importance weights in Eq. 2 are omitted for clarity and detailed in the supplementary material.